The K.E. of an electron in 3rd orbit of H like species is 6.04eV. The area of 3rd orbit for such species is??

Grade:12th pass

1 Answers

Vikas TU
7 years ago
The Kinetic energy is given.
Energy = 13.6z^2/9 = 6.04
we get z = 2.
Radii = 0.5209 * 9/2 = > 2.34405 A.
Area = pi*2.34405*2.34405 in Angstrom^2
would be the apprroximate area.
(approximating to circlar orbit.)